The mean of a population is 75 and the standard deviation is 14. The shape of the population is unknown. Determine the probability of each of the following occurring from this population.

a. A random sample of size 33 yielding a sample mean of 79 or more
b. A random sample of size 140 yielding a sample mean of between 73 and 77
c. A random sample of size 218 yielding a sample mean of less than 75.7

(Round all the values of z to 2 decimal places and final answers to 4 decimal places.)
Answer #1

Solution :

Given that,

mean = = 75

standard deviation = = 14

a )n = 33

= 75

= / n = 14 33 = 2.4371

a ) P ( >79 )

= 1 - P ( < 79)

= 1 - P ( - /) < (79 - 75 / 2.437 )

= 1 - P( z < 4 / 2.437 )

= 1 - P ( z < 1.64 )   

Using z table

= 1 - 0.9495

= 0.0505

Probability = 0.0505

b ) n = 140

= 75

= / n = 14 140 = 1.1832

d ) P (n 73 < < 77 )

P (n 73 - 75 / 1.1832) < ( - / ) < ( 77 - 75 / 1.1832)

P ( - 2./ 1.1832 < z < 2 / 1.1832 )

P (-1.69 < z < 1.69)

P ( z < 1.69) - P ( z < -1.69)

Using z table

= 0.9545 - 0.0455

= 0.9090

Probability = 0.9090

c ) n = 218

= 75

= / n = 14218 =0.9482

P (   < 75.7 )

P ( - /) < (75.7 - 75 / 0.9482)

P ( z < 0.7/ 0.9482)

P ( z < 0.74 )

Using z table

= 0.7704

Probability = 0.7704